three fourths of the distance from 5 to 13
Answer by stanbon(75887)

three fourths of the distance from 5 to 13
----
(3/4)(13-5) = (3/4)(8) = 6
----------------------
If you are looking for the number that is
(3/4) the distance it is 5+6 = 11
=====================
Cheers,
Stan H.